Suppose the exact charge neutrality does not hold in a world and the electron has a charge 1%less in magnitude than the proton. Calculate the coulomb force acting between two blocks of iron each of mass 1kg separated by a distance of 1m. Number of protons in an iron atom=26 and 58kg of iron contains 6*10^26 atoms.
1
Expert's answer
2019-07-30T09:14:50-0400

Electric force between two charged object are:


"F=\frac{kQ_1Q_2}{r^2}"


Now, we need to find out the charge of the iron samples:

Each bar has a mass of 1 kg.

So, the number of atoms in each bar is: "N_a=m\/n=6*10^{26}\/58=1.0344*10^{25} atoms"

Each atom has 26 protons: "N_p=26N_a=2.69*10^{26} protons"


The exceeding charge of the each rod is: "Q=0.01N_a*q_e=4.304*10^5 C"


Finally, the force between two rods:


"F=\frac{kQ_1Q_2}{r^2}=\frac{9*10^9 *4.304^2*10^{10}}{1^2}=1.46*10^{21}N"